OpenStudy (anonymous):

let f(x)=x^3-4. find f(a+3)

OpenStudy (anonymous):

whre you see 'x' , change it with 'a+3'

OpenStudy (anonymous):

\[f(a+3)=(a+3)^3-4\]

OpenStudy (mathteacher1729):

\[f(\spadesuit)=(\spadesuit)^3-4\] \[f(x)=(x)^3-4\] \[f(\Box)=(\Box)^2-4\] \[f(a+3)=(a+3)^2-4\] etc.

OpenStudy (anonymous):

incorrect with square :P

OpenStudy (mathteacher1729):

Oh poo. Did I square it instead of cube it? :( The darn font is so tiny.

OpenStudy (anonymous):

next time instead of square use cube :D

OpenStudy (mathteacher1729):

I'll just set my zoom trigger to 200%. (Right click math symbols, settings, zoom trigger..) ;-p

OpenStudy (anonymous):

but with zoom factor you will see it only when question is posted, not while editing or not?

OpenStudy (anonymous):

Please help! classify this conic equations: 1) \[x ^{2}+y ^{2}+2x+2y+1=0\] 2) \[2x^{2}+2y^{2}-4x+4y+19=0\] 3) \[x^{2}+4y^{2}=100\] 4) \[8x ^{2}+3y ^{2}=120\] 5) \[y=x^{2}-2x+3\]

OpenStudy (anonymous):

so is that the final answer or can i simplify that?

OpenStudy (anonymous):

you can

OpenStudy (cornzyblack):

that is f(a + 3)= (a + 3)^{3} - 4 = a^3 +3a^2 + 9a + 27 - 4 = a^3 + 3a^2 + 9a + 23

OpenStudy (anonymous):

oh ok iwow this is tough lol